OpenStudy (anonymous):

Earth takes a year to orbit the sun at a distance of 1.5x10^11m. Calculate its speed.

OpenStudy (amistre64):

1 cycle per year ;)

OpenStudy (amistre64):

... 1 rpy?

OpenStudy (anonymous):

First calculate the circumference: D=2πr=2*(3.14)*1.5*10^11=9.42*10^11

OpenStudy (anonymous):

nah ..i want the speed ?? in km/s or m/s

OpenStudy (anonymous):

Then divide D with time (1 year) and find the speed.

OpenStudy (anonymous):

One year in seconds is 3.15*10^7 so the speed in m/s is: 9.42*10^11/3.15*10^7=0.30*10^5 m/s!

OpenStudy (anonymous):

thanku !! the answer is correct !! but still i didn't understand :'(

OpenStudy (anonymous):

The speed is given by dividing the distance with time. The distance the Earth travels is the circumference around the sun. Then you convert the year in seconds and you do the math!

OpenStudy (anonymous):

thanku !! the answer is correct !! but still i didn't understand :'(

OpenStudy (anonymous):

how did u take 1.5x10^11m as radius ??